OpenStudy (anonymous):

Write an equation of the ellipse with the given characteristics and center at (0,0) Vertex: (3,0) Co-vertex: (0,-2)

OpenStudy (mertsj):

If the vertex is (3,0), then a = 3 and x^2=9 If the co-vertex is (0,-2) then b = 2 and b^2=4 So the equation must be: \[\frac{x^2}{9}+\frac{y^2}{4}=1\]